Sapting Learning While hauling a log in the back of a flatbed truck are pulled over by the state police. Although the cant roll sideways, the police claim that the log could have slid out the back of the truck when accelerating from rest. You claim that the truck couldnt possibly accelerate at the level needed to achieve such an effect. Regardless, the police write a ticket anyway and now your day in court is The log has a mass of m-824 kg; the truck has a mass of M = 9.10 x 103 kg. According to the truck manufacturer, the truck can accelerate from 0 to 55 mph in 27.0 seconds, but this does not account for the additional mass of the log. Calculate the minimum coefficient of static friction needed to keep the log in the back of the truck Number to the scientific literature, the coefficient of static friction between the log and the trailer bed should y 0.840. Given this fact, answer the following be r There is little chance the log could have slid out the back. The police officer is correct. The log posed a real danger O The results are not conclusive Previous Check Answer Next Exit Hint
Answer #1

a = acceleration of the truck

Vi = initial velocity = 0 m/s

Vf = final velocity = 55 mph = 24.6 m/s

t = time = 27 sec

acceleration is given as

a = (Vf - Vi)/t

a = (24.6 - 0)/27

a = 0.91 m/s2

for the log not to move

fs = static frictional force = ma

us mg = ma

us = a/g = 0.91/9.8 = 0.093

since the 0.840 > us

hence there is little chance the log could have slid out the back